Ran across this question. Any help would be great. If you could can you please show your work. Thanks
Let f(t)=2-3t^2 Find
(A)f(-2)(B)f(-t)(c)-f(t) (D)-f(-t)

Answer by algebrahouse.com(1659)

f(t) = 2 - 3t²

A.) f(-2)
= 2 - 3(-2)² {substituted -2 for t}
= 2 - 3(4) {evaluated exponent}
= 2 - 12 {multiplied}
= -10 {subtracted}

B.) f(-t)
= 2 - 3(-t)² {substituted -t for t}
= 2 - 3t² {evaluated exponent}

C.) -f(t)
= -(2 - 3t²) {took negative of entire right side}
= -2 + 3t² {distributed negative sign}

D.) -f(-t)
f(-t) was 2 - 3t² {see problem B.}
= -(2 - 3t²) {took negative of entire right side}
= -2 + 3t²

For more help from me, visit: www.algebrahouse.com